support answer.  All cash flows are end-of-period unless otherwise stated .

5. You plan on retiring forty years after you begin your first job and wish to have $2,500,000 at retirement. You intend to make your first payment at the end of your first year on the job and then increase your payments by 3 percent annually thereafter. If you invest in a fund paying an annual percentage rate (APR) of six percent, compounded semiannually, what is the amount of the first payment necessary to reach your objective ?

6. Fast forward: You have just retired and reached your objective (see above) of accumulating $2,500,000 in your retirement fund. You expect to live for another thirty years (exactly) and wish to leave $1,000,000 to Illinois Tech when you are gone. What is the most that you could withdraw from this fund at the end of each month and still leave the money to Tech? The fund is expected to return eight per annum, compounded monthly?
Answer #1

5. Formula:

  

Therefore,

2,500,000 = P [ (1+0.0609)^40 - (1+0.03)^40 ] / (0.0609 - 0.03)

P = 10,470.37392

6.

Using BA II Plus calculator:

N = 12* 30 = 360 ; I/Y = 8/12 = 0.67; PV = - 2,500,000 ; FV = 1,000,000; PMT = ?

PMT = 17,673.135